# AI-Assisted Human Evaluation of Machine Translation

### v1

@@ 4.2 ESAAI for Evaluation of WMT Systems
Our goal is for ESAAI to be as reliable or more than ESA in ranking MT systems. We consider MQM collected by Freitag et al. (2023) as the human gold standard and show the system-level correlations with our protocol in App. Figure 9. Both ESA and ESAAI have similar correlations with MQMWMT, justifying our setup. In Table 2 we show that this protocol does not stray far away from existing ones in terms of segment-level rating. Many of these cross-protocol correlations are on part with inter-annotator agreement, which is naturally the upper bound. Notably, ESAAI has higher correlation than ESA or MQM by Kocmi et al. (2024) alone.
ESA ESAAI MQM GEMBA
@@
Table 2: Kendall $\tau$ segment-level correlations between evaluation protocols, ESA and ESAAI use direct scores.
@@
GEMBA is recall-focused and therefore the occurrence of “false positive” segments is low. In 89% of cases, spans that were marked by GEMBA to have 0 errors retained 0 errors after annotation (App. Table 4), and such segments had an average score of 95. This makes it possible to also use GEMBA as a pre-filtering step. If we replace all such segments with 100 (to not overfit), all but one system comparisons remain the same. Alternatively, one can also filter segments for which GEMBA marks 0 errors for most systems, which has the advantage that we do not alter the data. For this method, again all but one system comparison would be the same (Figure 4). Pre-filtering can thus result in almost 25% budget saving (~52 segments per system).
Figure 4: Average system scores with either substitution or filtering of segments with no GEMBA errors.
@@
With large enough evaluation, even noisy annotation schemes yield the true system ordering. Conversely, only robust annotation schemes are good on a small scale. We formalize this in Appendix A to show that ESAAI leads to better annotations than ESA or MQM. For each subset size, e.g. 30 source sentences, we select 1000 random subsets and compute the system ranking accuracy against the whole dataset. Results in Figure 5 show that GEMBA alone is the most consistent because of the lack of inter-annotator confusion. However, it also increases the stability and quality of scores that annotators assign manually in ESAAI. In addition in practice, one can annotate fewer examples (e.g. 2000 for ESAAI) to obtain the same system-level accuracy as a slower protocol (e.g. 2500 for ESA), lowering the costs.
Figure 5: Accuracy of a system ranking only on a subset against ranking on full data. Percentages are averages. See numbers in App. Table 8.
@@ 5 Conclusion
Our AI-assisted protocol of human evaluation of MT is faster and cheaper. This protocol is more robust and self-consistent and increases inter-annotator agreement by priming the annotators with pre-annotated error spans. Our analysis also shows that the annotators did not over-rely on the AI and were able to maintain evaluation quality. The inclusion of AI in evaluation also opens many options for further evaluation economy.
@@
Lastly, we use GEMBA, a GPT4-based system, for the quality estimation and work with WMT 2023 data. Unfortunately, we can not exclude the possibility of the QE system being trained on this data, though the texts and scores are kept in two separate large files with non-linear mappings.
@@ Appendix A Subset Consistency Formalization
Appendix B Learning to Annotate
Kocmi et al. (2024) showed that despite ESA being faster than MQM, the users learn to perform the MQM annotations slightly faster. We show similar results in Figure 7, though in our case the workers learn to perform the post-editing of GEMBA error span annotation slightly faster (0.18s faster with every segment). This effect is present despite the ESA annotators being at an advantage because there was fewer of them and they thus each individually processed more segments. Even though the speedup happens thorough the whole annotation, it is mostly present in the first 15% of segments (green box in Figure 7). For ESA, this is -2.1s/segment and for ESAAI this is -1.9s/segment. In addition, users in the post-editing task seem to be more consistent. For ESA, the user’s deviation from their personal average is 43.3s, while for post-editing GEMBA this is only 32.1s. Overall, this makes the human effort more consistent and predictable but also showcases that the nature of the annotation task changes.
We now examine what specifically makes some segments take longer than others. We do so using feature-level correlations as shown in Table 3. Many of these features are co-dependent. For example, the longer the translation, the more likely GEMBA finds more error spans and the lower the final score. Nevertheless it gives us insights that ESAAI users learn to become faster. The number of words in the translation, together with the number of error spans is a strong predictor of annotation time. For MQM this is the highest, which can be explained by each error span requiring the most work in the MQM annotation scheme because the annotators have to also assign the error type. The longer the document (number of translation paragraphs), the lower the annotation time, which is likely due to shared context. With longer documents, the annotator does not have to switch between domains and contexts. Contrastively, the ESAAI annotators are slightly less affected by the translation length in contrast to ESA.
MQM ESA ESAAI
Progress -0.12 -0.13 -0.13
Translation word count 0.30 0.19 0.16
GEMBA error spans 0.12 0.07 0.12
Error spans 0.06 0.04 0.12
Score -0.07 -0.03 -0.08
Document size -0.14 -0.17 -0.17
Appendix C User Guidelines
@@
GEMBA ——Removed—— No edit ——Added——
#err. (freq.) $=$2 $=$1 $=$0 $=$0 $=$1 $=$2 $\geq$3
0 (23.8%) 0% 0% 100% 88% 88% 8% 2% 2%
1 (38.0%) 0% 28% 72% 62% 81% 14% 3% 3%
2 (18.8%) 15% 16% 69% 54% 71% 13% 9% 7%
3 (10.4%) 11% 20% 62% 51% 68% 16% 7% 10%
4 (8.9%) 11% 13% 69% 54% 65% 13% 10% 12%
Table 4: Distribution of error span post-editing based on original GEMBA-reported error spans (2nd column). Percentages in the table are proportions within the number of GEMBA error spans. For example, second row shows that 62% of segments with exactly one GEMBA error span received no post-editing from annotators and in 28% the annotators removed the single error. ESA is comparable to ESAAI.
Operation Frequency
Severity change 12.0%
                Increase severity 60.0%
                Decrease severity 40.0%
Move span $\leq$5 13.1%
Move span $\leq$10 17.2%
Move span $\leq$20 23.3%
Resize      Increase error span size 21.5%
                Decrease error span size 78.5%
Table 5: Distribution of two ESAAI post-editing types: changing the severity, and moving the error span. A span is considered to be moved if the distance between old and new endpoints is at most 5, 10, or 20 characters. Many GEMBA errors are only misplaced or have the wrong severity. See specific cases in Table 9.
SRC:   Sie haben gestern das Treffen wieder verschoben.
TGT:   He postponed the meeting again yesterday.
[…]
{example}
[…]
Original Perturbed OK
ESA Score 79.5 52.6 86%
Span count 0.85 1.86 54%
Perturbation marked 56%
ESAAI Score 75.8 52.6 76%
Span count 2.19 4.48 61%
Perturbation marked 71%
Scoring ESA ESAAI
direct score 0.376 0.533
from spans 0.327 0.671
Table 7: Inter-annotator agreement with direct scores and scores computed from error spans with MQM formula, as measured with Spearman correlation. ESAAI spans have the highest inter-annotator agreement, which is however caused by the GEMBA pre-filling. Still, the scores from ESAAI, solely by humans, have the highest inter-annotator agreement.
Protocol/ method Subset size
10 40 115 190
ESAAI 84.41% 92.38% 96.69% 98.88%
[…]
GEMBAspans 85.73% 93.10% 96.86% 98.94%
ESA 81.86% 90.26% 95.52% 98.52%
ESAspans 78.11% 88.28% 94.48% 97.94%
MQMspans 77.19% 86.30% 93.89% 98.50%
Table 8: Specific values of Figure 5. Subset accuracy across annotation schemes. ESAAI spans has the highest subset consistency, though this is likely biased by the spans from GEMBA, which as 100% inter-annotator agreement. However, ESAAI (direct scores) is based solely on human scorings, which has the second-highest subset consistency of any protocol.
@@
Figure 10: Changes in scoring by the same annotator when evaluated again. Each point represents single annotated segment with x-axis being annotator’s score assigned in March and y-axis their score assigned in May. ESAAI has the highest intra-annotator agreement, showing another positive aspect of being primed by GEMBA.
GEMBA Die Physik ist schrecklich und die Leute, die das Spiel entwickelt haben, werden nichts dagegen tun
Decrease severity Source Will not buy Mr. Coffee again
GEMBA Ich kaufe Mr. Kaffee nicht mehr.
ESAAI Ich kaufe Mr. Kaffee nicht mehr.
GEMBA Aber ich hasse Kunst und Literatur, und meine Schulgeschichte bestätigt es.
Resize Source [$\ldots$] I’m not sure if that would work for this.
GEMBA [$\ldots$] ich bin mir nicht sicher, ob das für diesen Zweck funktionieren würde.
{example}

### v2

@@ 4.2 ESAAI for Evaluation of WMT Systems
Our goal is for ESAAI to be as reliable or more than ESA in ranking MT systems. We consider MQMWMT collected by Freitag et al. (2023) as the human gold standard and show the system-level correlations with our protocol in App. Figure 9. Both ESA and ESAAI have similar correlations with MQMWMT, justifying our setup. In Table 7 we show that this protocol does not stray far away from existing ones in terms of segment-level rating. Many of these cross-protocol correlations are on part with inter-annotator agreement, which is naturally the upper bound. Notably, ESAAI has higher correlation than ESA or MQM by Kocmi et al. (2024) alone.
ESA ESAAI MQM QE
@@
Table 7: Kendall $\tau_{c}$ segment-level correlations between evaluation protocols, ESA and ESAAI use direct scores.
@@
The goal is to speed-up the annotation process without sacrificing quality. This can be achieved by removing, or automating, redundant decisions and actions on the annotator’s side. In this segment, we do so by skipping high-quality segments for which we can predict that the annotator’s would not mark any error spans.
Our QE system, GEMBA is recall-focused and therefore the occurrence of “false positive” error spans is low. In 89% of cases, spans that were marked by the QE to have 0 errors retained 0 errors after annotation (Table 8), and such segments had an average score of 95. This makes it possible to also use the QE as a pre-filtering step. If we replace all such segments with 100 (to not overfit), all but one system comparisons remain the same. Alternatively, one can also filter segments for which the QE marks 0 errors for most systems, which has the advantage that we do not alter the data. For this method, again all but one system comparison would be the same (Figure 7). Pre-filtering can thus result in almost 25% budget saving (~52 segments per system).
QE ——Removed—— No edit ——Added——
#err. (freq.) $=$2 $=$1 $=$0 $=$0 $=$1 $=$2 $\geq$3
0 (23.8%) 0% 0% 100% 88% 88% 8% 2% 2%
1 (38.0%) 0% 28% 72% 62% 81% 14% 3% 3%
2 (18.8%) 15% 16% 69% 54% 71% 13% 9% 7%
3 (10.4%) 11% 20% 62% 51% 68% 16% 7% 10%
4 (8.9%) 11% 13% 69% 54% 65% 13% 10% 12%
Table 8: Distribution of error span post-editing based on original QE-reported error spans (2nd column). Percentages in the table are proportions within the number of the QE error spans. For example, second row shows that 62% of segments with exactly one QE error span received no post-editing from annotators and in 28% the annotators removed the single error. ESA is comparable to ESAAI.
Figure 7: Average system scores with either substitution or filtering of segments with no QE errors. Each cross is a single system. The two red pairs of crosses are the single pairs of systems whose ordering changes when substitution or filtering is applied.
@@
Comparing the quality of two annotation protocols is not straightforward because of the absence of a gold standard. We now take a practical perspective where we desire a protocol that finds the true system ordering with as few examples as possible. This does not require any gold standard to compare to. Instead, this approach compares a subset of the annotations of a protocol to the full set of annotations from the same protocol.
Figure 8: Subset consistency accuracy of a system ranking only on a subset against ranking on full data. Percentages are averages across visible subset sizes, which corresponds to the area under the curve. See figure in tabular form in App. Table 9.
With large enough evaluation, even noisy annotation schemes yield the true system ordering. Conversely, only robust annotation schemes yield this ordering on a small scale. We formalize this in Appendix A to show that ESAAI leads to better annotations than ESA or MQM. We measure the accuracy of the ordering ($m_{1}{>_{I}}m_{2}$) of systems ($\mathcal{M}$) computed on a subset of segments ($I$) against the ordering given by the full data ($a_{m_{1}}{>}a_{m_{2}}$):
[…]
This simulates the setup where we wish to save costs with fewer number of annotations but care about arriving at the true system ordering. We refer to this metric as subset consistency accuracy and in Appendix A justify that annotation protocols with lower noise reach higher accuracies. For each subset size, e.g. 30 source sentences, we select 1000 random subsets and compute the system ranking accuracy.
Results in Figure 8 suggest that the QE system, GEMBA, alone is the most consistent. However, this is solely because it can be perceived as a single annotator and thus there is no inter-annotator confusion. From human evaluation protocols, ESAAI with scores has the highest stability and quality of scores. This is not the result of automation bias because only the annotators themselves assign the final score. In practice this would mean that one can annotate fewer examples (e.g. 2000 for ESAAI) to obtain the same system-level accuracy as lower-quality protocol (e.g. 2500 for ESA), thus further lowering the costs.
@@ 5 Conclusion
Our AI-assisted protocol of human evaluation of MT is faster and cheaper. This protocol is more robust and self-consistent and increases inter-annotator agreement by priming the annotators with pre-annotated error spans. Our analysis also shows that the annotators did not over-rely on the AI and were able to maintain evaluation quality. The inclusion of AI in evaluation also opens many options for further evaluation economy by reducing the test set size requirements. To this end, we introduce subset consistency accuracy, which quantifies how many annotations could be saved while arriving at similar final system ordering.
@@
Lastly, for QE we use GEMBA, a GPT4-based system, for the quality estimation and work with WMT 2023 data. Unfortunately, we can not exclude the possibility of the QE system being trained on this data, though the texts and scores are kept in two separate large files with non-linear mappings.
@@ Appendix A Subset Consistency Formalization
Appendix B User Guidelines
@@
Figure 10: Inter-annotator and intra-annotator agreement. The-intra annotator agreement shows changes in scoring by the same annotator when evaluated again. Each point represents single annotated segment with x-axis being annotator’s score assigned one month and y-axis their score assigned two months later. For ESA and ESAAI, the scores are directly from annotators. For MQM, they are computed by the formula. ESAAI has the highest intra-annotator and inter-annotator agreement, showing another positive aspect of being primed by GEMBA.
@@
Appendix C GEMBA Quality Estimator Prompts
Your task is to identify machine translation errors and assess the tranlsation quality.
{source_lang} source:
‘‘‘{source_seg}‘‘‘
{target_lang} translation:
‘‘‘{target_seg}‘‘‘
Based on the source segment and machine translation surrounded with triple backticks,
identify error types in the translation and classify them. The categories of errors are:
accuracy (addition, mistranslation, omission, untranslated text), fluency (character enc-
oding, grammar, inconsistency, punctuation, register, spelling), style (awkward), termin-
ology (inappropriate for context, inconsistent use), non-translation, other, or no-error.
Each error is classified as one of two categories: major or minor. Major errors disrupt
the flow and make the understandability of text difficult or impossible. Minor errors
are errors that do not disrupt the flow significantly and what the text is trying to say
is still understandable.
(a) Prompt for annotating error spans (initial step).
"source_lang": "English",
"source_seg": "I do apologise about this, we must gain permission from the account holder to discuss an order with another person, I apologise if this was done previously, however, I would not be able to discuss this with yourself without the account holders permission.",
"target_lang": "German",
"target_seg": "Ich entschuldige mich dafuer, wir muessen die Erlaubnis einholen, um eine Bestellung mit einer anderen Person zu besprechen. Ich entschuldige mich, falls dies zuvor geschehen waere, aber ohne die Erlaubnis des Kontoinhabers waere ich nicht in der Lage, dies mit dir involvement.",
"answer: """\
Major:
accuracy/mistranslation - "involvement"
accuracy/omission - "the account holder"
Minor:
fluency/grammar - "waere"
fluency/register - "dir"
(b) Prompt for scoring with prior annotations of error spans (example in prompt).
Given the translation from {source_lang} to {target_lang} and the annotated error
spans, assign a score on a continuous scale from 0 to 100. The scale has following
reference points: 0="No meaning preserved", 33="Some meaning preserved", 66="Most
meaning preserved and few grammar mistakes", up to 100="Perfect meaning and grammar".
Score the following translation from {source_lang} source:
‘‘‘{source_seg}‘‘‘
{target_lang} translation:
‘‘‘{target_seg}‘‘‘
Annotated error spans:
‘‘‘{error_spans}‘‘‘
Score (0-100):
(c) Prompt for scoring with prior annotations of error spans (final step).
Figure 12: Prompts for GEMBA with GPT-4. See the full GEMBA code for ESA. The prompts can be used to first prompt GEMBA to produce the list of translation errors, as in MQM, and then prompt again to score the segments hollistically. For the ESAAI human pre-annotations we only use the first part and only the error severities.